# Anchoring and Mixed-Norm Contractions in Averaging-Learning Dynamics

### v1

@@ 2.3 The case where AA-$\mathcal{E}$ can contain negative diagonal entries
Let $A$ be a row-stochastic matrix of size $n>0$ and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$. Suppose that $0\leq\mathcal{E}(i)\leq 2A(i,i)$ for all $i\in\{1,2,\dots,n\}$. If AA-$\mathcal{E}$ is condensely anchored then AA-$\mathcal{E}$ is zero-convergent.
@@
This implies that each row-sum of $\lvert A-\mathcal{E}\rvert$ is at most $1$, and hence, $\lvert A-\mathcal{E}\rvert$ is row substochastic. Then, there exists a row-stochastic matrix $\bar{A}$ and a diagonal nonnegative matrix $\bar{\mathcal{E}}$ such that $\lvert A-\mathcal{E}\rvert=\bar{A}-\bar{\mathcal{E}}\geq 0$. If $i\in\{1,2,\dots,n\}$ is an anchor of AA-$\mathcal{E}$, then
@@
where $\lvert A(i,i)-\mathcal{E}(i)\rvert<A(i,i)$ means that the $i$-th row-sum of $\lvert A-\mathcal{E}\rvert=\bar{A}$-$\bar{\mathcal{E}}$ is less than $1$. By Proposition 2.4, we deduce that $i\in\{1,2,\dots,n\}$ is an anchor of $\bar{A}$-$\bar{\mathcal{E}}$ if and only if $i$ is an anchor of AA-$\mathcal{E}$. The digraphs $G[A-\mathcal{E}]$ and $G[\bar{A}-\bar{\mathcal{E}}]$ are also identical. Hence, since AA-$\mathcal{E}$ is condensely anchored, then $\bar{A}$-$\bar{\mathcal{E}}$ is also condensely anchored. By Theorem 2.11, we have that $\bar{A}$-$\bar{\mathcal{E}}$ is zero-convergent. By Theorem 1.1, we have $\rho(\bar{A}-\bar{\mathcal{E}})<1$. Since
@@
then AA-$\mathcal{E}$ is zero-convergent, also by Theorem 1.1. ∎
@@
Let $A$ be a row-stochastic matrix of size $n>0$ and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$. Suppose that $0\leq\mathcal{E}(i)\leq 2A(i,i)$ for all $i\in\{1,2,\dots,n\}$. If AA-$\mathcal{E}$ is condensely anchored then all agents eventually reach consensus, i.e., $\lim_{t\to\infty}X_{t}=\bar{\sigma}$.
@@
Note that $A-\mathcal{E}-I_{n}/n$ has rank $2$, i.e., it has eigenvalue 0 of multiplicity n2n-2. Let $\lambda_{1}$ and $\lambda_{2}$ be the remaining two eigenvalues of $A-\mathcal{E}-I_{n}/n$. Since $\mathrm{tr}(A-\mathcal{E}-I_{n}/n)=-1$ and $\mathrm{tr}((A-\mathcal{E}-I_{n}/n)^{2})=1+2/n-2/n^{3}$, we have
@@
Let $A$ be a row-stochastic matrix of size $n\geq 2$ and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$. Let $A(1,1)=1$ and for all $i\in\{2,3,\dots,n\}$, let A(i,i1)=1A(i,i-1)=1. Then, AA-$\mathcal{E}$ is zero-convergent if and only if $0<\mathcal{E}(1)<2$ and $0\leq\mathcal{E}(i)<1$ for all $i\in\{2,3,\dots,n\}$. This is clear since
@@
Let $A$ be a row-stochastic matrix of size $n>0$ and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$. Suppose that there exists an $i\in\{1,2,\dots,n\}$ such that $\mathcal{E}(i)\geq 3$. Is it true that $\rho(A-\mathcal{E})\geq 1$?
@@ 2.4 Impaired averaging
While the Friedkin-Johnsen model produces substochastic dynamics through intentional anchoring to initial opinions, we now consider a fundamentally different form of substochasticity arising from unintentional dissipation. Formally, if some rows of the averaging matrix $A$ do not sum to one, the system becomes proper substochastic. This corresponds to a loss of mass in the update, as though information is dissipating at each step. Concrete interpretations include: An agent discounts available signals so that $\sum_{j=1}^{n}A_{t}(i,j)<1$ for some $i$. An agent forgets part of its state when forming $X_{t+1}$ from $X_{t}$. And finally, external dissipation, opinion strength leaks out of the system. In the absence of stubborn agents or exogenous learning inputs, such impairments force the state toward the zero vector, which becomes the unique fixed point of the dynamics. This contrasts sharply with FJ, where the missing weight is preserved through anchoring to $X_{0}$, yielding persistent disagreement at a nontrivial equilibrium. In faulty averaging, the missing weight simply vanishes: consensus through extinction rather than compromise. In this sense, impaired averaging can substitute for perturbations or feedback control: if a row-sum is less than $1$, then the spectral radius drops below $1$, which ensures asymptotic decay. The following corollary of Theorem 2.11 makes this precise.
@@
Let $A$ be a proper-substochastic matrix of positive size. Then $A$ is zero-convergent if and only if $A$ is condensely anchored.
@@ 3.1 Vanishing learning rates
In the earlier regime $0<\mathcal{E}_{t}(i)<2A_{t}(i,i)$ but now we add the possibility that learning ability is vanishing. Each agent eventually has positive self-belief, however, no additional structure is required on the interaction matrix $A$. Condition (A3) ($\sum_{t=0}^{\infty}\min_{i}\mathcal{E}_{t}(i)=\infty$) is analogous to the slow cooling condition in simulated annealing. As established by [24, 30], this condition ensures sufficient ’total cooling’ to overcome any finite energy barrier and reach the global optimum. In our context, with effective temperature $T_{t}=1/\mathcal{E}_{\min,t}$ condition (A3) guarantees that $\int_{0}^{\infty}\mathcal{E}_{\min,t}dt=\infty$, allowing the system to escape any metastable state (wrong consensus) and reach consensus.
@@ 5 Conclusion
In this paper, we have shown that the Averaging-Learning dynamics allow for quite nuanced behaviour. Agents can be social but not learn through the learning-rate matrix $\mathcal{E}$. This, however, imposes conditions on the averaging matrix $A$. The first half of the paper developed the concept of anchoring, using graph theory. The condensed anchoring property is not only new but useful to analyze networks where not all agents have access to perfect learning. In this sense, even when some agents are defective, society as a whole can converge to the truth.
@@
Several questions remain open. Problem 2.20 asks whether a learning rate of $3$ for any agent forces instability. A positive answer would give a sharp threshold for overlearning. The gap between necessity and sufficiency for the zero-convergence of AA-$\mathcal{E}$ when AA-$\mathcal{E}$ has negative diagonal entries also deserves further study.
@@ Appendix A Appendix: The Open Problem
Is it true that $\rho(A-\mathcal{E})\geq 1$ for $\mathcal{E}$ an $n\times n$ diagonal matrix such that for some $i\in\{1,2\dots,n\}$, $\mathcal{E}(i)\geq 3$ and $A$ is any $n\times n$ row-stochastic matrix?
@@
In the following theorem, we have that if $A$ is a symmetric row-stochastic matrix with an agent of learning rate at least $2$, then AA-$\mathcal{E}$ is not zero-convergent.
@@
Let $A$ be a symmetric row-stochastic matrix of size $n>0$ and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$. Suppose that there exists an $i\in\{1,2,\dots,n\}$ such that $\mathcal{E}(i)\geq 2$. Then $\rho(A-\mathcal{E})\geq 1$.
@@
If $\lambda$ is an eigenvalue of $A$ then $\lambda\leq 1$. On the other hand, if $\lambda$ is the smallest eigenvalue of -$\mathcal{E}$ then $\lambda\leq-\mathcal{E}(i)\leq-2$. By the Weyl’s theorem [32, Theorem 4.3.1], if $\lambda$ is the smallest eigenvalue of AA-$\mathcal{E}$ then $\lambda\leq-2+1=-1$, which implies that $\rho(A-\mathcal{E})\geq\lvert\lambda\rvert\geq 1$. ∎
@@
If the matrix $A$ in the assumption of Theorem A.2 is not symmetric, then it is possible that $\rho(A-\mathcal{E})<1$. For example, let
@@
Let $A$ be a row-stochastic matrix of size $n>0$ and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$. Suppose that there exists an $i\in\{1,2,\dots,n\}$ such that $\mathcal{E}(i)\geq 2$. Suppose also that $0\leq\mathcal{E}(j)<2A(j,j)$ for all $j\in\{1,2,\dots,n\}$ where $j\neq i$. Then $\rho(A-\mathcal{E})\geq 1$.
@@ A.1 The quantitative version
If $A$ is row-stochastic and $\mathcal{E}$ is diagonal, then lower bounds on $f_{n}(R)$ translate into lower bounds for $\rho(A-\mathcal{E})$ when $\max_{i}\mathcal{E}(i)=R$.
@@ A.2.2 Application to $f_{n}(R)$ and a threshold $R_{n}$ for $f_{n}(R)>1$
where $\mathcal{E}=\mathrm{diag}(\mathcal{E}(1),\dots,\mathcal{E}(n))$ with $\max_{i}\mathcal{E}(i)=\mathcal{E}(1)=R$ and $A$ is substochastic (row sums $\leq 1$).
@@
Because $A$ is substochastic, $\|A\|_{\infty}\leq 1$. Moreover $\|N\|_{\infty}=\|\mathcal{E}\|_{\infty}=R$ and one checks row-wise that

### v2

@@ 2.3 The case where AA-$\mathcal{E}$ can contain negative diagonal entries
Let AA be a row-stochastic matrix of size $n>0$ and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$. Suppose that $0\leq\mathcal{E}(i)\leq 2A(i,i)$ for all $i\in\{1,2,\dots,n\}$. If AA-$\mathcal{E}$ is condensely anchored then AA-$\mathcal{E}$ is zero-convergent.
@@
This implies that each row-sum of $\lvert A-\mathcal{E}\rvert$ is at most 11, and hence, $\lvert A-\mathcal{E}\rvert$ is row substochastic. Then, there exists a row-stochastic matrix $\bar{A}$ and a diagonal nonnegative matrix $\bar{\mathcal{E}}$ such that $\lvert A-\mathcal{E}\rvert=\bar{A}-\bar{\mathcal{E}}\geq 0$. If $i\in\{1,2,\dots,n\}$ is an anchor of AA-$\mathcal{E}$, then
@@
where $\lvert A(i,i)-\mathcal{E}(i)\rvert<A(i,i)$ means that the $i$-th row-sum of $\lvert A-\mathcal{E}\rvert=\bar{A}$-$\bar{\mathcal{E}}$ is less than 11. By Proposition 2.4, we deduce that $i\in\{1,2,\dots,n\}$ is an anchor of $\bar{A}$-$\bar{\mathcal{E}}$ if and only if $i$ is an anchor of AA-$\mathcal{E}$. The digraphs $G[A-\mathcal{E}]$ and $G[\bar{A}-\bar{\mathcal{E}}]$ are also identical. Hence, since AA-$\mathcal{E}$ is condensely anchored, then $\bar{A}$-$\bar{\mathcal{E}}$ is also condensely anchored. By Theorem 2.11, we have that $\bar{A}$-$\bar{\mathcal{E}}$ is zero-convergent. By Theorem 1.1, we have $\rho(\bar{A}-\bar{\mathcal{E}})<1$. Since
@@
then AA-$\mathcal{E}$ is zero-convergent, also by Theorem $1.$ ∎
@@
Let AA be a row-stochastic matrix of size $n>0$ and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$. Suppose that $0\leq\mathcal{E}(i)\leq 2A(i,i)$ for all $i\in\{1,2,\dots,n\}$. If AA-$\mathcal{E}$ is condensely anchored then all agents eventually reach consensus, i.e., $\lim_{t\to\infty}X_{t}=\bar{\sigma}$.
@@
Note that $A-\mathcal{E}-I_{n}/n$ has rank 22, i.e., it has eigenvalue 0 of multiplicity n2n-2. Let $\lambda_{1}$ and $\lambda_{2}$ be the remaining two eigenvalues of $A-\mathcal{E}-I_{n}/n$. Since $\mathrm{tr}(A-\mathcal{E}-I_{n}/n)=-1$ and $\mathrm{tr}((A-\mathcal{E}-I_{n}/n)^{2})=1+2/n-2/n^{3}$, we have
@@
Let AA be a row-stochastic matrix of size $n\geq 2$ and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$. Let $A(1,1)=1$ and for all $i\in\{2,3,\dots,n\}$, let A(i,i1)=1A(i,i-1)=1. Then, AA-$\mathcal{E}$ is zero-convergent if and only if $0<\mathcal{E}(1)<2$ and $0\leq\mathcal{E}(i)<1$ for all $i\in\{2,3,\dots,n\}$. This is clear since
@@
Let AA be a row-stochastic matrix of size $n>0$ and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$. Suppose that there exists an $i\in\{1,2,\dots,n\}$ such that $\mathcal{E}(i)\geq 3$. Is it true that $\rho(A-\mathcal{E})\geq 1$?
@@ 2.4 Impaired averaging
While the Friedkin-Johnsen model produces substochastic dynamics through intentional anchoring to initial opinions, we now consider a fundamentally different form of substochasticity arising from unintentional dissipation. Formally, if some rows of the averaging matrix AA do not sum to one, the system becomes proper substochastic. This corresponds to a loss of mass in the update, as though information is dissipating at each step. Concrete interpretations include: An agent discounts available signals so that $\sum_{j=1}^{n}A_{t}(i,j)<1$ for some ii. An agent forgets part of its state when forming $X_{t+1}$ from $X_{t}$. And finally, external dissipation, opinion strength leaks out of the system. In the absence of stubborn agents or exogenous learning inputs, such impairments force the state toward the zero vector, which becomes the unique fixed point of the dynamics. This contrasts sharply with FJ, where the missing weight is preserved through anchoring to $X_{0}$, yielding persistent disagreement at a nontrivial equilibrium. In faulty averaging, the missing weight simply vanishes: consensus through extinction rather than compromise. In this sense, impaired averaging can substitute for perturbations or feedback control: if a row-sum is less than 11, then the spectral radius drops below 11, which ensures asymptotic decay. The following corollary of Theorem 2.11 makes this precise.
@@
Let AA be a proper-substochastic matrix of positive size. Then AA is zero-convergent if and only if AA is condensely anchored.
@@ 3.1 Vanishing learning rates
In the earlier regime $0<\mathcal{E}_{t}(i)<2A_{t}(i,i)$ but now we add the possibility that learning ability is vanishing. Each agent eventually has positive self-belief, however, no additional structure is required on the interaction matrix AA. Condition (A3) ($\sum_{t=0}^{\infty}\min_{i}\mathcal{E}_{t}(i)=\infty$) is analogous to the slow cooling condition in simulated annealing. As established by [24, 30], this condition ensures sufficient ’total cooling’ to overcome any finite energy barrier and reach the global optimum. In our context, with effective temperature $T_{t}=1/\mathcal{E}_{\min,t}$ condition (A3) guarantees that $\int_{0}^{\infty}\mathcal{E}_{\min,t}dt=\infty$, allowing the system to escape any metastable state (wrong consensus) and reach consensus.
@@ 5 Conclusion
In this paper, we have shown that the Averaging-Learning dynamics allow for quite nuanced behaviour. Agents can be social but not learn through the learning-rate matrix $\mathcal{E}$. This, however, imposes conditions on the averaging matrix AA. The first half of the paper developed the concept of anchoring, using graph theory. The condensed anchoring property is not only new but useful to analyze networks where not all agents have access to perfect learning. In this sense, even when some agents are defective, society as a whole can converge to the truth.
@@
Several questions remain open. Problem 2.20 asks whether a learning rate of 33 for any agent forces instability. A positive answer would give a sharp threshold for overlearning. The gap between necessity and sufficiency for the zero-convergence of AA-$\mathcal{E}$ when AA-$\mathcal{E}$ has negative diagonal entries also deserves further study.
@@ Appendix A Appendix: The Open Problem
Is it true that $\rho(A-\mathcal{E})\geq 1$ for $\mathcal{E}$ an $n\times n$ diagonal matrix such that for some $i\in\{1,2\dots,n\}$, $\mathcal{E}(i)\geq 3$ and AA is any $n\times n$ row-stochastic matrix?
@@
In the following theorem, we have that if AA is a symmetric row-stochastic matrix with an agent of learning rate at least 22, then AA-$\mathcal{E}$ is not zero-convergent.
@@
Let AA be a symmetric row-stochastic matrix of size $n>0$ and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$. Suppose that there exists an $i\in\{1,2,\dots,n\}$ such that $\mathcal{E}(i)\geq 2$. Then $\rho(A-\mathcal{E})\geq 1$.
@@
If $\lambda$ is an eigenvalue of AA then $\lambda\leq 1$. On the other hand, if $\lambda$ is the smallest eigenvalue of -$\mathcal{E}$ then $\lambda\leq-\mathcal{E}(i)\leq-2$. By the Weyl’s theorem [32, Theorem 4.3.1], if $\lambda$ is the smallest eigenvalue of AA-$\mathcal{E}$ then $\lambda\leq-2+1=-1$, which implies that $\rho(A-\mathcal{E})\geq\lvert\lambda\rvert\geq 1$. ∎
@@
If the matrix AA in the assumption of Theorem A.2 is not symmetric, then it is possible that $\rho(A-\mathcal{E})<1$. For example, let
@@
Let AA be a row-stochastic matrix of size $n>0$ and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$. Suppose that there exists an $i\in\{1,2,\dots,n\}$ such that $\mathcal{E}(i)\geq 2$. Suppose also that $0\leq\mathcal{E}(j)<2A(j,j)$ for all $j\in\{1,2,\dots,n\}$ where $j\neq i$. Then $\rho(A-\mathcal{E})\geq 1$.
@@ A.1 The quantitative version
If AA is row-stochastic and $\mathcal{E}$ is diagonal, then lower bounds on $f_{n}(R)$ translate into lower bounds for $\rho(A-\mathcal{E})$ when $\max_{i}\mathcal{E}(i)=R$.
@@ A.2.2 Application to $f_{n}(R)$ and a threshold $R_{n}$ for $f_{n}(R)>1$
where $\mathcal{E}=\mathrm{diag}(\mathcal{E}(1),\dots,\mathcal{E}(n))$ with $\max_{i}\mathcal{E}(i)=\mathcal{E}(1)=R$ and AA is substochastic (row sums $\leq 1$).
@@
Because AA is substochastic, $\|A\|_{\infty}\leq 1$. Moreover $\|N\|_{\infty}=\|\mathcal{E}\|_{\infty}=R$ and one checks row-wise that